Solve each system by the addition method.
3/7x+5/9y=27
1/9x+2/7y=7
and
3x-2.5y=7.125
2.5x-3y=7.3125

Answer by Alan3354(69443)

Solve each system by the addition method.
3/7x+5/9y=27
1/9x+2/7y=7
Since it says "linear systems" I'll assume 3/7x means (3/7)x
(3x/7) + (5y/9) = 27 Do them like this to avoid ambiguities.
(x/9) + (2y/7) = 7
Multiply both eqns by 63 to eliminate fractions.
27x + 35y = 1701
7x + 18y = 441
Multiply the 1st eqn by 7 and the 2nd by 27 to get 189x in both
189x + 245y = 11907
189x + 486y = 11907
Subtract eqn 2 from eqn 1
0x - 241y = 0
y = 0
x = 11907/189
x = 63
-------------
and
3x-2.5y=7.125
2.5x-3y=7.3125
--------------
Multiply both by 16
48x - 40y = 114
40x - 48y = 117
All integers now
Multiply the 1st eqn by 5, the 2nd by 6
240x - 200y = 570
240x - 288y = 702
Subtract the 2nd from the 1st
0x + 88y = -132
y = -1.5
-------------
Sub for y into 2.5x-3y=7.3125
2.5x -3*(-1.5) = 7.3125
2.5x + 4.5 = 7.3125
2.5x = 2.8125
x = 1.125